# Ctrl-V: Higher Fidelity Video Generation with Bounding-Box Controlled Object Motion

### v1

@@ Abstract
With recent advances in video prediction, controllable video generation has been attracting more attention. Generating high fidelity videos according to simple and flexible conditioning is of particular interest. To this end, we propose a controllable video generation model using pixel level renderings of 2D or 3D bounding boxes as conditioning. In addition, we also create a bounding box predictor that, given the initial and ending frames’ bounding boxes, can predict up to 15 bounding boxes per frame for all the frames in a 25-frame clip. We perform experiments across 3 well-known AV video datasets: KITTI, Virtual-KITTI 2 and BDD100k. Project page: https://oooolga.github.io/ctrl-v.github.io/.
@@ 2 Related Work
Motion Control with Bounding-Box Conditioning There are many strategies of control that have been explored in controllable video generation research. Notably, ControlVideo [22] utilizes a training-free strategy that employs pre-trained image LDMs and ControlNets to generate videos based on canny edge and depth maps. Control-A-Video [19] leverages a controllable video LDM that combines a pre-trained text-to-video model with ControlNet to manipulate videos using similar visual information. Video ControlNets [20, 21] uses optical flow information to enhance video generation, while ReVideo [23] depends on extracted video trajectories. DreamPose [18] injects pose sequence information into the initial noise. VideoComposer [15] uses an array of sketch, depth, mask, and motion vectors as conditioning. Many of these conditions, such as edge, depth, and optical flow maps, are costly to produce and lack the flexibility needed for customization. Bounding box emerge as a conditioning that are easily customizable and can be edited into different shape, size, locations and classes efficiently. To the best of our knowledge, five other research projects are currently exploring the use of bounding boxes for motion control in video generation. However, it is important to note that our work is distinct from these in several critical respects. TrailBlazer [24] and Peekaboo [25] are different training-free approaches that employ attention map adjustments to direct the model in generating a particular object within a defined region. FACTOR [26] augmented the transformer-based generation model, Phenaki [27], by integrating a box control module. TrailBlazer, Peekaboo and FACTOR necessitate textual descriptions for individual boxes, thus lacking direct visual grounding.
@@ 3.2 Preprocessing: Bounding-Box Frame Preparation
Instead of resorting to Fourier transform or other sequential models to encode bounding-box information in string format, we have chosen to represent the bounding boxes through image plotting. We created a method to make bounding-box plots which have significant benefits. These plots not only show the precise locations of objects within a scene but also effectively display different kinds of information, like track IDs and object types, using different shapes and colors. We also implement a method to render trajectory frames, similar to the bounding-box frames, but without the pixel-space size information of the objects. These trajectory frames only contain mid-point location information. The procedure for creating bounding-box/trajectory plots is detailed in Appendix A.1. Additionally, we utilize the image-encoder ($\mathcal{E}$) to encoder the bounding-box frames, and its performance is discussed in Appendix A.2.
@@ 4 Experimental Analysis and Ablation Studies
Furthermore, each bounding-box predictor discussed in the following section requires conditioning on ONE initial frame, THREE initial bounding-box frames and ONE final bounding-box/trajectory frame.
@@ 4.2 Generation Quality
4.3 Bounding-Box Predictor: Quantitative Evaluation
@@ C.2 BDD Result Visualizations: Bounding-Box Predictions and Motion-Controlled Video Generations
In the following section, we showcase a range of BDD generation results produced by our model in different scene scenarios, such as city, urban, highways, busy intersections and at night. Each visualization displays every 6th frame from a 25-frame clip, with the actual video generated at a frame rate of 5 fps. In the leftmost column labels, GT represents ground truth, PB represents predicted bounding-box frames, and PF represents predicted frames.
@@
In the following section, we showcase a range of vKITTI generation results produced by our model in different scene scenarios. In Figure 14, we showcase our generations on the vKITTI test split. However, it is worth noting that the vKITTI test split comprises samples from only one type of scene (unseen during training). To provide a broader persepective, in Figure 15, we present additional results on the vKITTI train split. Each visualization displays every 6th frame from a 25-frame clip, with the actual video generated at a frame rate of 7 fps. In the leftmost column labels, GT represents ground truth, PB represents predicted bounding-box frames, and PF represents predicted frames.
@@ C.4 KITTI Result Visualizations: Bounding-Box Predictions and Motion-Controlled Video Generations
In the following section, we showcase a range of KITTI generation results produced by our model in different scene scenarios, such as urban area, city streets and highway. Each visualization displays every 6th frame from a 25-frame clip, with the actual video generated at a frame rate of 7 fps. In the leftmost column labels, GT represents ground truth, PB represents predicted bounding-box frames, and PF represents predicted frames.
@@ D.1 Noteworthy Discoveries in our Box2Video Model
Our experiments have yielded findings suggesting that our Box2Video model not only understands bounding-box locations in scenes but also encodes track IDs and 3D bounding-box orientation information. While these findings are not conclusive and require further investigation to be confirmed, they highlight potential directions for future projects. Here we share some of these findings to guide further research and development.
@@
    Given the previous findings, we believe that drawing our bounding-box outlines a bit thicker to make them more prominent could potentially improve our results.

### v2

@@ Abstract
With recent advances in video prediction, controllable video generation has been attracting more attention. Generating high fidelity videos according to simple and flexible conditioning is of particular interest. To this end, we propose a controllable video generation model using pixel level renderings of 2D or 3D bounding boxes as conditioning. In addition, we also create a bounding box predictor that, given the initial and ending frames’ bounding boxes, can predict up to 15 bounding boxes per frame for all the frames in a 25-frame clip. We perform experiments across 3 well-known AV video datasets: KITTI, Virtual-KITTI 2 and BDD100k.
Project page: https://oooolga.github.io/ctrl-v.github.io/.
@@ 2 Related Work
Motion Control with Bounding-Box Conditioning There are many strategies of control that have been explored in controllable video generation research. Notably, ControlVideo [22] utilizes a training-free strategy that employs pre-trained image LDMs and ControlNets to generate videos based on canny edge and depth maps. Control-A-Video [19] leverages a controllable video LDM that combines a pre-trained text-to-video model with ControlNet to manipulate videos using similar visual information. Video ControlNets [20, 21] uses optical flow information to enhance video generation, while ReVideo [23] depends on extracted video trajectories. DreamPose [18] injects pose sequence information into the initial noise. VideoComposer [15] uses an array of sketch, depth, mask, and motion vectors as conditioning.
Many of these conditions, such as edge, depth, and optical flow maps, are costly to produce and lack the flexibility needed for customization. Bounding box emerge as a conditioning that are easily customizable and can be edited into different shape, size, locations and classes efficiently. To the best of our knowledge, five other research projects are currently exploring the use of bounding boxes for motion control in video generation. However, it is important to note that our work is distinct from these in several critical respects.
TrailBlazer [24] and Peekaboo [25] are different training-free approaches that employ attention map adjustments to direct the model in generating a particular object within a defined region. FACTOR [26] augmented the transformer-based generation model, Phenaki [27], by integrating a box control module. TrailBlazer, Peekaboo and FACTOR necessitate textual descriptions for individual boxes, thus lacking direct visual grounding.
@@ 3.2 Preprocessing: Bounding-Box Frame Preparation
Instead of resorting to Fourier transform or other sequential models to encode bounding-box information in string format, we have chosen to represent the bounding boxes through image rendering. We created a method to make bounding-box plots which have significant benefits. These plots not only show the precise locations of objects within a scene but also effectively display different kinds of information, like track IDs and object types, using different shapes and colors. We also implement a method to render trajectory frames, similar to the bounding-box frames, but without the pixel-space size information of the objects. These trajectory frames only contain mid-point location information. The procedure for creating bounding-box/trajectory plots is detailed in Appendix A.1. Additionally, we utilize the image-encoder ($\mathcal{E}$) to encoder the bounding-box frames, and its performance is discussed in Appendix A.2.
@@ 4 Experimental Analysis and Ablation Studies
Furthermore, each bounding-box predictor discussed in the following section requires conditioning on ONE initial frame, ONE or THREE initial bounding-box frames and ONE final bounding-box/trajectory frame.
@@ 4.2 Generation Quality
BBox Predictor + Box2Video 1-to-1 467.7 0.3416 0.3241 13.21
BBox Predictor + Box2Video 3-to-1 422.2 0.2981 0.4277 13.85
@@
BBox Predictor + Box2Video 1-to-1 400.2 0.3179 0.4714 15.78
BBox Predictor + Box2Video 3-to-1 341.4 0.2645 0.5841 17.60
@@
BBox Predictor + Box2Video 1-to-1 412.8 0.2967 0.5470 17.52
BBox Predictor + Box2Video 3-to-1 373.1 0.3071 0.5407 17.37
[…]
@@
4.3 BBox Predictor: Quantitative Evaluation
# Cond.
Frames
@@ 5 Conclusions
Acknowledgments
We thank CIFAR for their support under the AI chairs program, NSERC for their support under the Discovery Grants program, and Samsung for supporting this work.
@@ C.2 BDD Result Visualizations: Bounding-Box Predictions and Motion-Controlled Video Generations
In the following section, we showcase a range of BDD generation results produced by our 3-to-1 generation pipeline in different scene scenarios, such as city, urban, highways, busy intersections and at night. Each visualization displays every 6th frame from a 25-frame clip, with the actual video generated at a frame rate of 5 fps. In the leftmost column labels, GT represents ground truth, PB represents predicted bounding-box frames, and PF represents predicted frames.
@@
In the following section, we showcase a range of vKITTI generation results produced by our 3-to-1 generation pipeline in different scene scenarios. In Figure 14, we showcase our generations on the vKITTI test split. However, it is worth noting that the vKITTI test split comprises samples from only one type of scene (unseen during training). To provide a broader persepective, in Figure 15, we present additional results on the vKITTI train split. Each visualization displays every 6th frame from a 25-frame clip, with the actual video generated at a frame rate of 7 fps. In the leftmost column labels, GT represents ground truth, PB represents predicted bounding-box frames, and PF represents predicted frames.
@@ C.4 KITTI Result Visualizations: Bounding-Box Predictions and Motion-Controlled Video Generations
In the following section, we showcase a range of KITTI generation results produced by our 3-to-1 generation pipeline in different scene scenarios, such as urban area, city streets and highway. Each visualization displays every 6th frame from a 25-frame clip, with the actual video generated at a frame rate of 7 fps. In the leftmost column labels, GT represents ground truth, PB represents predicted bounding-box frames, and PF represents predicted frames.
@@ D.1 Noteworthy Discoveries in our Box2Video Model
Our experiments have yielded findings suggesting that our Box2Video model not only understands bounding-box locations in scenes but also the encoded track IDs and 3D bounding-box orientation information in the rendered bounding-box frames. While these findings are not conclusive and require further investigation to be confirmed, they highlight potential directions for future projects. Here we share some of these findings to guide further research and development.
@@
    Given the previous findings, we believe that plotting the bounding-box outlines a bit thicker to make them more prominent could potentially improve our results.